# Neutrino mixing parameters and masses from $\Delta(96)\rtimes H_{CP}$ in the tri-direct CP approach

### v1

@@ 4.4 The breaking pattern $\mathcal{T}_{10}$
The lepton mixing angles and the Jarlskog invariant can be naturally derived from the PMNS matrix. By comprehensively scanning over the parameter space of $m_{a}$, $x$, $r$ and $\eta$, the predicted ranges for input parameters, three lepton mixing angles, Dirac CP phase, Majorana CP phase and neutrino masses are presented in table LABEL:tab:viable_regions. Furthermore, we obtain the best fit values of the input and output parameters, as shown in table 3.
@@
which leads to the following the sum rule among the Dirac CP phase $\delta_{CP}$ and mixing angles
@@
In this case, the output parameters such as lepton mixing angles and the absolute mass masses are all solely determined by two input parameters $m_{a}$ and $r$. By comprehensively scanning over the parameter space of $m_{a}$ and $r$, the corresponding input parameters, neutrino masses and mixing parameters are constrained to the ranges below:
@@
Note that eliminating the input parameter $r$, we can relate mixing parameters and mass ratios to the reactor mixing angle $\sin\theta_{13}$, and we present them in the figure 8.
@@ 4.5 The breaking pattern $\mathcal{S}_{2}$
The neutrino mass matrix $m^{\prime}_{\nu}$ can be further diagonalized by the unitary transformation $U_{\nu 2}$. Thus, PMNS matrix can be obtained in the following form via the systematic procedure described in Eq. (2.21)
@@
The explicit expressions for the mixing parameters can then be easily derived. We present the best fit values of neutrino mixing parameters and neutrino masses with input parameters $m_{a}$$x$, $r$ and $\eta$ all being free parameters, as listed in table 3. The specific ranges of various parameters can be obtained and the results are presented in table LABEL:tab:viable_regions.
@@
Figure 9: Contour plot of $\delta_{CP}/\pi$ within $\sin^{2}\theta_{13}-\sin^{2}\theta_{23}$ plane, and the dimensionless observable quantities $\sin^{2}\theta_{23}$, $\sin^{2}\theta_{13}$, $\delta_{CP}$ and $m_{2}^{2}/m_{3}^{2}$ in the $\eta$/$\pi-r$ plane for the breaking pattern $\mathcal{S}_{2}$ with $x=(1+\sqrt{11})/2$.
As an illustrative example, we choose the solar alignment parameter $x=(1+\sqrt{11})/2$ which gives a relatively simple values of the first column of the PMNS matrix
@@
Furthermore, we can derive the following sum rules among the mixing angles $\theta_{12}$ and $\theta_{13}$
@@
The left panel of figure 9 illustrates the contour map of $\delta_{CP}/\pi$ in the $\sin^{2}\theta_{13}-\sin^{2}\theta_{23}$ plane. The red region represents the resulting area obtained by scanning parameters $m_{a}$$\eta$ and $r$ within a reasonable range, followed by constraining all observable quantities according to their $3\sigma$ experimental ranges. From the figure, it is evident that the Dirac CP phase $\delta_{CP}$ is concentrated around $0.9\pi$. Moreover, we plot the contour regions for the $3\sigma$ experimental ranges of mixing angles $\theta_{13}$, $\theta_{23}$, the Dirac CP phase $\delta_{CP}$ and mass ratio $m_{2}/m_{3}$ in the plane $\eta$ and $r$ in right panel of figure 9.
The numerical results for lepton mixing angles and neutrino masses arising from two additional special $\eta$ are tabulated in table 5. In the following analysis, the phase parameter $\eta$ is constrained to $\eta=4\pi/5$, a value that can be easily realized in a concrete model. All lepton mixing parameters and the neutrino mass ratio $m_{2}/m_{3}$ are now determined by a single dimensionless parameter $r$. This leads to the predictions for mixing parameters and neutrino masses interesting. Figure 10 illustrates the functional correlation of lepton mixing parameters, the Dirac CP phase $\delta_{CP}$, the Majorana CP phase $\beta$ and mass ratio $m_{2}/m_{3}$ on $\sin\theta_{13}$. Upon further scanning of parameters $m_{a}$ and $r$, it is found that the input parameters and output parameters are constrained to
@@ 5 Conclusion
The type I seesaw mechanism elegantly explains the exceptionally smallness of neutrino masses on a theoretical level, yet its experimental verification remains challenging. The Majorana masses of RHNs could be extremely heavy, which may lie beyond the reach of the most advanced particle colliders. Furthermore, the seesaw mechanism introduces more free parameters than measurable quantities and is hard to test experimentally. In the present work, guided by the principles of minimality and symmetry, we explore a highly constrained and predictive variant of the seesaw model: the minimal seesaw model with two right-handed neutrino based on the tri-direct CP approach. In this kind of models, the neutrino masses, mixing angles and CP violation phases are determined from just three or four real input parameters.
@@
We emphasise that the model independent tri-direct CP approach is a quite predictive scheme for constructing neutrino mass models based on discrete flavor symmetry and CP symmetry. In tri-direct CP approach, we have analyzed all possible breaking patterns which arise from the high energy symmetry $\Delta(96)\rtimes H_{CP}$ based on the two right-handed neutrino seesaw mechanism. The symmetry $\Delta(96)\rtimes H_{CP}$ is spontaneously broken down to an Abelian subgroup $G_{l}$ (non $Z_{2}$ subgroups) in the charged lepton sector, to $G_{\text{atm}}\rtimes H^{\text{atm}}_{CP}$ in the atmospheric neutrino sector and to $G_{\text{sol}}\rtimes H^{\text{sol}}_{CP}$ in the solar neutrino sector. In this work, we assign the left-handed lepton doublets $L$ transform as the $\Delta(96)$ triplet $\bm{3_{0}}$. The flavon field $\phi_{\text{atm}}$, coupling to $N_{\text{atm}}$ and $L$, transforms as the conjugate triplet $\bm{\bar{3}_{0}}$. The flavon $\phi_{\text{sol}}$, coupling to $N_{\text{sol}}$ and $L$, transforms as either $\bm{\bar{3}_{0}}$ or $\bm{\bar{3}_{1}}$. For the former assignment, the neutrino masses, lepton mixing angles and CP violation phases of each breaking pattern are determined by three real input parameters $|m_{a}|$, $r$ and $\eta$. After performing an exhaustive numerical analysis for all possible independent breaking patterns, we find out only one phenomenologically interesting mixing pattern which agree with the experimental data very well. The corresponding representative combination of residual flavor symmetries is $\{G_{l},G_{\text{atm}},G_{\text{sol}}\}=(Z_{3}^{ac},Z_{2}^{a^{2}b},Z_{3}^{ad^{2}})$, and both the atmospheric and solar neutrino sectors exhibit trivial residual CP transformations represented by the identity matrix $\mathbb{1}_{3}$. The residual symmetries lead to the vacuum alignments $\langle\phi_{\text{atm}}\rangle=(0,1,-1)^{T}v_{\phi_{a}}$ and $\langle\phi_{\text{sol}}\rangle=(1,-1,1)^{T}v_{\phi_{s}}$. As a consequence, the lepton mixing matrix is predicted to be TM1 pattern, the neutrino masses are NO and the lightest neutrino is massless with $m_{1}=0$
@@
When the flavon $\phi_{\text{sol}}$ is considered to transform as $\bm{\bar{3}_{1}}$, a new real parameter $x$ must be included. We have analyzed all possible symmetry breaking patterns in a model independent way, including the corresponding predictions for lepton mixing parameters and neutrino masses with NO mass hierarchy. We find 42 phenomenologically interesting mixing patterns with $\chi^{2}_{\text{min}}$ less than 10: 27 feature the charged lepton residual symmetry $G_{l}$ as $Z^{ac}_{3}$ (labeled $\mathcal{T}_{i}$), 7 exhibit $Z^{c}_{4}$ (labeled $\mathcal{F}_{i}$), 5 demonstrate $Z^{abd^{2}}_{4}$ (labeled $\mathcal{S}_{i}$), and 3 display $Z^{abd}_{8}$ (labeled $\mathcal{W}_{i}$). The corresponding combination of residual flavor symmetries of each viable breaking pattern are summarized in table 2. The best fit values of input parameters, mixing parameters and neutrino masses are presented in table 3. Furthermore, after a comprehensive scan of the parameter space, the predicted ranges of input and output parameters of the 42 breaking patterns are derived, as shown in table LABEL:tab:viable_regions.
@@
Guided by the analysis of viable breaking patterns, we present detailed numerical results of the five cases $\mathcal{F}_{6}$, $\mathcal{F}_{7}$, $\mathcal{F}_{4}$, $\mathcal{T}_{10}$ and $\mathcal{S}_{2}$ for some benchmark values of $x$ and $\eta$. In the breaking pattern $\mathcal{F}_{6}$, the benchmark value $x=3$ leads to $\langle\phi_{\text{sol}}\rangle=(-1,i,3i)^{T}v_{\phi_{s}}$ and the so-called TM1 mixing matrix. For the breaking pattern $\mathcal{F}_{7}$, the benchmark values of $x$ are taken to be $x=3$ and $x=(2+\sqrt{17})/2\approx 3$. The former case is exactly the littlest seesaw model with CSD(3) which was originally proposed in Ref. [60]. For the latter case, the first column of PMNS matrix is predicted to be $\frac{1}{5}(\sqrt{17},2,2)^{T}$. The breaking pattern $\mathcal{F}_{4}$ shows the predictions for benchmark values $x=5$ and $\eta=5\pi/3$. In this case, the absolute value of the first column of PMNS matrix is fixed to be $\frac{1}{\sqrt{38}}(5,2,3)^{T}$. For the breaking pattern $\mathcal{T}_{10}$, to enhance the predictive power, we focus on $x=4$ and $\eta=\pi$, and derive that the corresponding first column of the PMNS matrix is $\frac{1}{\sqrt{57}}(\sqrt{37},\sqrt{10},\sqrt{10})^{T}$. We consider the case of $x=(1+\sqrt{11})/2$ in the breaking pattern $\mathcal{S}_{2}$. In the same way, the first column of the PMNS matrix is obtained as $\frac{1}{3}(\sqrt{6},1,\sqrt{2})^{T}$. While the TM1 mixing scheme is well established, the other four matrices presented here have not so far appeared in the literature. Our analysis reveals that these five distinct breaking patterns are highly effective in predicting lepton mixing parameters and neutrino mass values when $x$ and $\eta$ assume specific benchmark values. Crucially, all the lepton mixing parameters and the neutrino masses are restricted in rather narrow regions, and can easily be tested by the forthcoming neutrino experiments. In particular the predictions of solar mixing angles, atmospheric mixing angles and the normal neutrino mass hierarchy will be tested quite soon.

### v2

@@ 4.4 The breaking pattern $\mathcal{T}_{10}$
The lepton mixing angles and the Jarlskog invariant can be naturally derived from the PMNS matrix. By comprehensively scanning over the parameter space of $m_{a}$, $x$, $r$ and $\eta$, we present the predicted ranges for input parameters, three lepton mixing angles, Dirac CP phase, Majorana CP phase and neutrino masses in table LABEL:tab:viable_regions. Furthermore, we obtain the best fit values of the input and output parameters, as shown in table LABEL:tab:viable_best_fit.
@@
which leads to the following sum rule among the Dirac CP phase $\delta_{CP}$ and mixing angles
@@
In this case, the output parameters such as lepton mixing angles and the absolute masses are all solely determined by two input parameters $m_{a}$ and $r$. By comprehensively scanning over the parameter space of $m_{a}$ and $r$, we constrain the corresponding input parameters, neutrino masses and mixing parameters to the ranges below:
@@
After eliminating the input parameter $r$, we can relate mixing parameters and mass ratios to the reactor mixing angle $\sin\theta_{13}$, and we present them in the figure 9.
@@ 4.5 The breaking pattern $\mathcal{S}_{2}$
The neutrino mass matrix $m^{\prime}_{\nu}$ can be further diagonalized by the unitary transformation $U_{\nu 2}$. Thus, the PMNS matrix can be obtained in the following form via the systematic procedure described in Eq. (2.21)
@@
The explicit expressions for the mixing parameters can then be easily derived. We present the best fit values of neutrino mixing parameters and neutrino masses with input parameters $m_{a}$$x$, $r$ and $\eta$ all being free parameters, as listed in table LABEL:tab:viable_best_fit. The specific ranges of various parameters can be obtained and the results are presented in table LABEL:tab:viable_regions.
@@
Figure 10: Contour plot of $\delta_{CP}/\pi$ within $\sin^{2}\theta_{13}-\sin^{2}\theta_{23}$ plane, and the dimensionless observable quantities $\sin^{2}\theta_{23}$, $\sin^{2}\theta_{13}$, $\delta_{CP}$ and $m_{2}^{2}/m_{3}^{2}$ in the $\eta$$\pi-r$ plane for the breaking pattern $\mathcal{S}_{2}$ with $x=(1+\sqrt{11})/2$.
As an illustrative example, we choose the solar alignment parameter $x=(1+\sqrt{11})/2$ which gives relatively simple values of the first column of the PMNS matrix
@@
Furthermore, we can derive the following sum rule between the mixing angles $\theta_{12}$ and $\theta_{13}$
@@
The left panel of figure 10 illustrates the contour map of $\delta_{CP}/\pi$ in the $\sin^{2}\theta_{13}-\sin^{2}\theta_{23}$ plane. The red region represents the resulting area obtained by scanning parameters $m_{a}$$\eta$ and $r$ within a reasonable range, followed by constraining all observable quantities according to their $3\sigma$ experimental ranges. From the figure, it is evident that the Dirac CP phase $\delta_{CP}$ is concentrated around $0.9\pi$. Moreover, we plot the contour regions for the $3\sigma$ experimental ranges of mixing angles $\theta_{13}$, $\theta_{23}$, the Dirac CP phase $\delta_{CP}$ and mass ratio $m_{2}/m_{3}$ in the plane $\eta$ and $r$ in the right panel of figure 10.
The numerical results for lepton mixing angles and neutrino masses arising from two additional special $\eta$ are tabulated in table 5. In the following analysis, the phase parameter $\eta$ is constrained to $\eta=4\pi/5$, a value that can be easily realized in a concrete model. All lepton mixing parameters and the neutrino mass ratio $m_{2}/m_{3}$ are now determined by a single dimensionless parameter $r$. This leads to interesting predictions for mixing parameters and neutrino masses. Figure 11 illustrates the functional correlation of lepton mixing parameters, the Dirac CP phase $\delta_{CP}$, the Majorana CP phase $\beta$ and mass ratio $m_{2}/m_{3}$ on $\sin\theta_{13}$. Upon further scanning of parameters $m_{a}$ and $r$, it is found that the input parameters and output parameters are constrained to
@@
4.6 The breaking pattern $\mathcal{F}_{2}^{\prime}$
From table LABEL:tab:invariant_VEVs, we find that the VEVs of $\phi_{\text{atm}}$ and $\phi_{\text{sol}}$ are proportional to
$\langle\phi_{\text{atm}}\rangle=(1,-1,-1)^{T}v_{\phi_{a}},\qquad\langle\phi_{\text{sol}}\rangle=(1,i,ix)^{T}v_{\phi_{s}}\,.$ (4.46)
Consequently, the light neutrino mass matrix $m_{\nu}$ takes the form
$m_{\nu}=m_{a}\left[\begin{pmatrix}1&-1&-1\\ -1&1&1\\ -1&1&1\\ \end{pmatrix}+re^{i\eta}\begin{pmatrix}1&i&ix\\ i&-1&-x\\ ix&-x&-x^{2}\\ \end{pmatrix}\right$]\,, (4.47)
Using the form specified for the IO case in Eq. (2.18), $m_{\nu}$ can be transformed into a block-diagonal matrix $m^{\prime}_{\nu}$ via the unitary transformation $U_{\nu 1}$, where
$U_{\nu 1}=\begin{pmatrix}\frac{1}{\sqrt{3}}&-\frac{i(x+1+2i)}{\sqrt{6(x-1)x+12}}&-\frac{i(x-1)}{\sqrt{2(x-1)x+4}}\\ -\frac{1}{\sqrt{3}}&\frac{ix+1-2i}{\sqrt{6(x-1)x+12}}&\frac{-1-ix}{\sqrt{2(x-1)x+4}}\\ -\frac{1}{\sqrt{3}}&\frac{1+i-2ix}{\sqrt{6(x-1)x+12}}&\frac{1+i}{\sqrt{2(x-1)x+4}}\\ \end{pmatrix}\,$. (4.48)
The entries of $m^{\prime}_{\nu}$ are given by
$y=\frac{m_{a}}{3}\left[9-r(x+1+i)^{2}e^{i\eta}\right],~~z=\frac{m_{a}}{3}r(1-i-ix)\sqrt{2\left(x^{2}-x+2\right)}e^{i\eta},~~w=\frac{2m_{a}}{3}r(x^{2}-x+2)e^{i\eta}\,.$ (4.49)
Furthermore, $m^{\prime}_{\nu}$ can be diagonalized by the unitary matrix $U_{\nu 2}$ given in Eq. (2.20). Thus, the PMNS matrix can be written as
$U_{PMNS}=P_{321}U_{\nu 1}U_{\nu 2}\,.$ (4.50)
The explicit expressions for the mixing parameters can then be easily derived. When input parameters are treated as free parameters, we present the best fit values of the other output parameters in table LABEL:tab:viable_best_fit. By comprehensively scanning over space of input parameters, the phenomenologically viable ranges for $x$, $\eta$, $m_{a}$, $r$, three lepton mixing angles, Dirac CP phase, Majorana CP phase and neutrino masses are presented in table LABEL:tab:viable_regions.
For the benchmark value $x$ = 7, the absolute value of the third column of PMNS matrix is fixed to be:
$\frac{1}{2\sqrt{11}}\left(1,5,3\sqrt{2}\right)^{T},$ (4.51)
Then the expressions of the mixing angles and the Jarlskog invariant $J_{CP}$ can be read as
$\sin^{2}\theta_{13}=\frac{1}{44},\qquad\sin^{2}\theta_{12}=\frac{1}{258}\left[129-2\sqrt{22}\sin 2\theta(\cos\psi-13\sin\psi)-41\cos 2\theta\right],$
$\sin^{2}\theta_{23}=\frac{25}{43},\qquad J_{CP}=\frac{1}{968}\left[22\cos 2\theta-\sqrt{22}\sin\theta\cos\theta(4\sin\psi+11\cos\psi)\right]\,.$ (4.52)
Furthermore, we can derive the following sum rule between the mixing angles $\theta_{13}$ and $\theta_{23}$
$\cos^{2}\theta_{13}\cos^{2}\theta_{23}=\frac{9}{22}\,.$ (4.53)
In this case, the best fit values of the observables corresponding to different values of $\eta$ are summarized in table 5. In the following analysis, the phase parameter $\eta$ is constrained to $\eta=52\pi/27$. All lepton mixing parameters and the neutrino mass ratio are now determined by a single dimensionless parameter $r$. By combining the experimental data presented in table 1 and further scanning the parameters $m_{a}$ and $r$, we find that the observables are constrained to
$21.700\,\text{meV}\leq m_{a}\leq 21.898\,\text{meV},\qquad 0.0586\leq r\leq 0.0587\,$,
$\sin^{2}\theta_{13}=0.02273,\qquad 0.276\leq\sin^{2}\theta_{12}\leq 0.332,\qquad\sin^{2}\theta_{23}=0.581,$
$1.688\leq\delta_{CP}/\pi\leq 1.716,\qquad 1.622\leq\beta/\pi\leq 1.644,\qquad 13.801\,\text{meV}\leq m_{ee}\leq 16.544\,\text{meV},$
$m_{3}=0,\qquad 48.927\,\text{meV}\leq m_{1}\leq 49.432\,\text{meV},\qquad 49.643\,\text{meV}\leq m_{2}\leq 50.128\,\text{meV}\,.$ (4.54)
For the IO scenario, we can see that $\sin^{2}\theta_{13}$ and $\sin^{2}\theta_{23}$ are constant values.
@@ 5 Conclusion
The type I seesaw mechanism elegantly explains the exceptionally small of neutrino masses on a theoretical level, yet its experimental verification remains challenging. The Majorana masses of RHNs could be extremely heavy, which may lie beyond the reach of the most advanced particle colliders. Furthermore, the seesaw mechanism introduces more free parameters than measurable quantities and is hard to test experimentally. In the present work, guided by the principles of minimality and symmetry, we explore a highly constrained and predictive variant of the seesaw model: the minimal seesaw model with two right-handed neutrinos based on the tri-direct CP approach. In this kind of models, the neutrino masses, mixing angles and CP violation phases are determined from just three or four real input parameters.
@@
We emphasise that the model independent tri-direct CP approach is a quite predictive scheme for constructing neutrino mass models based on discrete flavor symmetry and CP symmetry. In the tri-direct CP approach, we have analyzed all possible breaking patterns which arise from the high energy symmetry $\Delta(96)\rtimes H_{CP}$ based on the two right-handed neutrino seesaw mechanism. The symmetry $\Delta(96)\rtimes H_{CP}$ is spontaneously broken down to an Abelian subgroup $G_{l}$ (non $Z_{2}$ subgroups) in the charged lepton sector, to $G_{\text{atm}}\rtimes H^{\text{atm}}_{CP}$ in the atmospheric neutrino sector and to $G_{\text{sol}}\rtimes H^{\text{sol}}_{CP}$ in the solar neutrino sector. In this work, we assign the left-handed lepton doublets $L$ to transform as the $\Delta(96)$ triplet $\bm{3_{0}}$. The flavon field $\phi_{\text{atm}}$, coupling to $N_{\text{atm}}$ and $L$, transforms as the conjugate triplet $\bm{\bar{3}_{0}}$. The flavon $\phi_{\text{sol}}$, coupling to $N_{\text{sol}}$ and $L$, transforms as either $\bm{\bar{3}_{0}}$ or $\bm{\bar{3}_{1}}$. For the former assignment, the neutrino masses, lepton mixing angles and CP violation phases of each breaking pattern are determined by three real input parameters $|m_{a}|$, $r$ and $\eta$. After performing an exhaustive numerical analysis for all possible independent breaking patterns, we find out only one phenomenologically interesting mixing pattern which agrees with the experimental data very well. The corresponding representative combination of residual flavor symmetries is $\{G_{l},G_{\text{atm}},G_{\text{sol}}\}=(Z_{3}^{ac},Z_{2}^{a^{2}b},Z_{3}^{ad^{2}})$, and both the atmospheric and solar neutrino sectors exhibit trivial residual CP transformations represented by the identity matrix $\mathbb{1}_{3}$. The residual symmetries lead to the vacuum alignments $\langle\phi_{\text{atm}}\rangle=(0,1,-1)^{T}v_{\phi_{a}}$ and $\langle\phi_{\text{sol}}\rangle=(1,-1,1)^{T}v_{\phi_{s}}$. As a consequence, the lepton mixing matrix is predicted to be the TM1 pattern, the neutrino masses are NO and the lightest neutrino is massless with $m_{1}=0$.
@@
When the flavon $\phi_{\text{sol}}$ is considered to transform as $\bm{\bar{3}_{1}}$, a new real parameter $x$ must be included. We have analyzed all possible symmetry breaking patterns in a model independent way, including the corresponding predictions for lepton mixing parameters and neutrino masses with both NO and IO mass hierarchy. We find 54 (42 for NO and 12 for IO) phenomenologically interesting mixing patterns with $\chi^{2}_{\text{min}}$ less than 10: 33 feature the charged lepton residual symmetry $G_{l}$ as $Z^{ac}_{3}$ (labeled $\mathcal{T}_{i}$ and $\mathcal{T}_{i}^{\prime}$), 9 exhibit $Z^{c}_{4}$ (labeled $\mathcal{F}_{i}$ and $\mathcal{F}_{i}^{\prime}$), 7 demonstrate $Z^{abd^{2}}_{4}$ (labeled $\mathcal{S}_{i}$ and $\mathcal{S}_{i}^{\prime}$), and 5 display $Z^{abd}_{8}$ (labeled $\mathcal{W}_{i}$ and $\mathcal{W}_{i}^{\prime}$). The corresponding combination of residual flavor symmetries of each viable breaking pattern is summarized in table 2. The best fit values of input parameters, mixing parameters and neutrino masses are presented in table LABEL:tab:viable_best_fit. Furthermore, after a comprehensive scan of the parameter space, the predicted ranges of input and output parameters of the 54 breaking patterns are derived, as shown in table LABEL:tab:viable_regions.
@@
Guided by the analysis of viable breaking patterns, we present detailed numerical results of the five NO cases $\mathcal{F}_{6}$, $\mathcal{F}_{7}$, $\mathcal{F}_{4}$, $\mathcal{T}_{10}$, $\mathcal{S}_{2}$ and one IO case $\mathcal{F}_{2}^{\prime}$ for some benchmark values of $x$ and $\eta$. In the breaking pattern $\mathcal{F}_{6}$, the benchmark value $x=3$ leads to $\langle\phi_{\text{sol}}\rangle=(-1,i,3i)^{T}v_{\phi_{s}}$ and the so-called TM1 mixing matrix. For the breaking pattern $\mathcal{F}_{7}$, the benchmark values of $x$ are taken to be $x=3$ and $x=(2+\sqrt{17})/2\approx 3$. The former case is exactly the littlest seesaw model with CSD(3) which was originally proposed in Ref. [60]. For the latter case, the first column of PMNS matrix is predicted to be $\frac{1}{5}(\sqrt{17},2,2)^{T}$. The breaking pattern $\mathcal{F}_{4}$ shows the predictions for benchmark values $x=5$ and $\eta=5\pi/3$. In this case, the absolute value of the first column of PMNS matrix is fixed to be $\frac{1}{\sqrt{38}}(5,2,3)^{T}$. For the breaking pattern $\mathcal{T}_{10}$, to enhance the predictive power, we focus on $x=4$ and $\eta=\pi$, and derive that the corresponding first column of the PMNS matrix is $\frac{1}{\sqrt{57}}(\sqrt{37},\sqrt{10},\sqrt{10})^{T}$. We consider the case of $x=(1+\sqrt{11})/2$ in the breaking pattern $\mathcal{S}_{2}$. In the same way, the first column of the PMNS matrix is obtained as $\frac{1}{3}(\sqrt{6},1,\sqrt{2})^{T}$. While the TM1 mixing scheme is well established, the other four matrices presented here have not so far appeared in the literature. For the breaking pattern $\mathcal{F}_{2}^{\prime}$, when $x$ = 7, the third column of the PMNS matrix corresponds to $\frac{1}{2\sqrt{11}}(1,5,3\sqrt{2})^{T}$. Our analysis reveals that these six distinct breaking patterns are highly effective in predicting lepton mixing parameters and neutrino mass values when $x$ and $\eta$ assume specific benchmark values. Crucially, all the lepton mixing parameters and the neutrino masses are restricted in rather narrow regions, and can easily be tested by the forthcoming neutrino experiments. In particular the predictions of solar mixing angles, atmospheric mixing angles and the normal neutrino mass hierarchy will be tested quite soon.